OpenStudy (anonymous):

A bacteria culture will grow from 400 to 5000 in 1 1/4 hours. a) Find the constant k for this bacteria using growth formula, y=ne^kt, and write the growth equation. b)Use your equation to predict how much longer it will take the 5000 bacteria to become 15000 bacteria.

OpenStudy (anonymous):

\[y=ne^{kt}\]

OpenStudy (anonymous):

n is beginning y is end

OpenStudy (anonymous):

Correct. Dont know where to start on this problem.

OpenStudy (anonymous):

so in beginning there was 400

OpenStudy (anonymous):

so plug in 400 for n

OpenStudy (anonymous):

plug in what for y?

OpenStudy (anonymous):

5000?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[5000=400 e^{kt}\]

OpenStudy (anonymous):

t represent time , what time do you see there

OpenStudy (anonymous):

t would go in as 1.5

OpenStudy (anonymous):

1 1/4= 1.25

OpenStudy (anonymous):

I knew that.

OpenStudy (anonymous):

let's go with that

OpenStudy (anonymous):

k. so the equation is \[5000=400e ^{k(1.25)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your job is to find k

OpenStudy (anonymous):

Hmm...take ln of both sides and bring exponent down then solve for k?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[\ln(5000)=k(1.25)\ln400e\]

OpenStudy (anonymous):

not what I would do

OpenStudy (anonymous):

How would you set it up?

OpenStudy (anonymous):

I would divide both sides by 400 first

OpenStudy (anonymous):

Ohh okay.

OpenStudy (anonymous):

\[5000/400=e ^{k(1.25)}\]

OpenStudy (anonymous):

now take ln of both sides

OpenStudy (anonymous):

\[\ln(25/2)=k(1.25)lne\]

OpenStudy (anonymous):

Wouldnt ln and e cancel eachother out though?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So basically, to get k, you would divide ln(25/2) by 1.25?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k=2.02?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol thanks. so to rewrite the equation, i would just plug everything in?

OpenStudy (anonymous):

\[5000=400e ^{2.02(1.25)}\]

OpenStudy (anonymous):

close but to be more general you use t instead of specific time

OpenStudy (anonymous):

Ohh okay. So intead of 1.25, just t?

OpenStudy (anonymous):

\[y=400e ^{2.02(t)}\]

OpenStudy (anonymous):

that way you can find y for any t value you plug in

OpenStudy (anonymous):

So on b, you would plug n as 5000 and y as 1500?

OpenStudy (anonymous):

*15000

OpenStudy (anonymous):

exactly , this time you are finding the 't'

OpenStudy (anonymous):

I got 0.5438

OpenStudy (anonymous):

Because you divide by 5000 on both sides. Getting: \[\ln(3)=(2.02)\ln(e)\]

OpenStudy (anonymous):

And then divide by ln(2.02) Did I do it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank God. lol. But would that be like 5 mins?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

.5438

OpenStudy (anonymous):

yes ,

OpenStudy (anonymous):

it will take .5438 hours

OpenStudy (anonymous):

okie dokes.

OpenStudy (anonymous):

Thanks youuuu!!

OpenStudy (anonymous):

nice to meet you

OpenStudy (anonymous):

Nice to meet you too. It was good working with you.